OpenStudy (anonymous):

what is tye inverse of f (x)=7x-1/2

OpenStudy (anonymous):

put f(x) = y \[y=7x-\frac{ 1 }{ 2 }\] then put \[x=7y-\frac{ 1 }{ 2 }\] \[7y=x+\frac{ 1 }{ 2 }\] \[y=\frac{ x+\frac{ 1 }{ 2 } }{ 7 }\] rewrite it as \[f(x)^{-1} \]\[f(x)^{-1} =\frac{ x+\frac{ 1 }{ 2 }}{ 7 }\]

OpenStudy (anonymous):

get it ??